OpenStudy (anonymous):

The price for a gallon of whole milk has increased from 2006 to 2008. From 2006 to 2007, the cost of whole milk increased by approximately 56%. From 2007 to 2008, the cost of milk whole milk increased by approximately 42%. If, on average, the cost for a gallon of whole milk nationwide was $1.99 at the beginning of 2006, determine the cost for a gallon of whole milk by the end of the year 2008. Round your answer to the nearest cent. a. $3.10 c. $2.83 b. $4.41 d. $3.94

OpenStudy (anonymous):

2006 was 1.99 per gallon From 2006 - 2007 cost increases 56 % : (1.99)(.56) = 1.12......1.99 + 1.12 = 3.11 From 2007 - 2008 cost increased 42% : (3.11)(.42) = 1.31 .....3.11 + 1.31 = 4.41 < -- answer